OpenStudy (kittiwitti1):

Find y if the point (15, y) is on the terminal side of θ and\[\cos\theta=\frac{15}{17}\](Enter your answers as a comma-separated list.) My result was \[\pm8\]but I'm told that it's the wrong answer. ッ

OpenStudy (jchick):

Can you post your work?

OpenStudy (kittiwitti1):

\[y=\sqrt{17^{2}-15^{2}}=\sqrt{289-225}=\sqrt{64}=\pm8\]

OpenStudy (kittiwitti1):

Okay, does that help or should I do anything else?

OpenStudy (jchick):

Did you do anything else?

OpenStudy (kittiwitti1):

\[x=15,y=8,r=17\]\[\sin\theta=\frac{8}{17},\cos\theta=\frac{15}{17},\csc\theta=\frac{17}{8},\sec\theta=\frac{17}{15}\]I would have kept going but I'm not sure if I'm going in the right direction.

OpenStudy (jchick):

Yes you are in the right direction that is what I got.

OpenStudy (kittiwitti1):

It is the wrong answer apparently...

OpenStudy (jchick):

Did you submit it?

OpenStudy (jchick):

That is odd

OpenStudy (kittiwitti1):

I put \[\pm8\]Wrong answer, it says.

OpenStudy (jchick):

Wait did you have to type your answer?

OpenStudy (kittiwitti1):

Yes. (I put -8, 8)

OpenStudy (jchick):

Does it automatically correct it?

OpenStudy (kittiwitti1):

@jim_thompson5910 @ganeshie8 @Hero No, it does not.

OpenStudy (kittiwitti1):

@DanJS @dan815

OpenStudy (jchick):

Ok so your teacher said it was wrong

OpenStudy (kittiwitti1):

No, the website says I am wrong.

OpenStudy (kittiwitti1):

http://prntscr.com/8di0sp

OpenStudy (jchick):

Ok that is a auto correcting website?

OpenStudy (jchick):

So maybe you put an extra space or something?

OpenStudy (kittiwitti1):

...No?

OpenStudy (jchick):

Ok sorry I can't help further

OpenStudy (kittiwitti1):

...Okay thanks anyway....

OpenStudy (kittiwitti1):

@DanJS can you help? :c

OpenStudy (kittiwitti1):

...I'm just going to repost this.